# Quantifying Classifier Utility under Local Differential Privacy

### v2

@@ 7.2.2 Bank Customer Attrition
Figure 7: Empirical and theoretical utility for two classifiers trained on the Bank Customer Attrition dataset.
Two key observations emerge in this case study: (i) The Exponential mechanism performs comparably to the PM mechanism for Logistic Regression. This is because its robustness hyperrectangle $\theta_{\diamond}=[0,0.72]\times[0,1]$ is large enough to cover most of the sensitive feature ranges, enabling excellent performance for all mechanisms defined on $[0,1]$. However, Laplace and Gaussian mechanisms are less effective here due to their probability mass extending beyond $[0,1]$. (ii) The Gaussian mechanism is not as good as the other mechanisms. The main reason is from the noise scale $\sigma$ in Theorem 4. Although we can extend the Gaussian mechanism to $\varepsilon\geq 1$, this leads to a larger $\sigma$, i.e. more noise added to the data.
@@
7.2.3 MNIST-$\times7$
Figure 8 illustrates the empirical and theoretical utility for a Neural Network classifier trained on the MNIST-$\times7$ dataset. We defer the detailed discussions, average-case and worst-case analyses to Appendix C.6.
@@ 7.2.3 MNIST-$\times7$
Figure 8: Empirical and theoretical utility for a Neural Network classifier trained on the MNIST-$\times7$ dataset.
7.3 Summary of Case Studies
The key findings from our case studies are as follows:
@@ 7.3 Summary of Case Studies
8 Conclusions
This paper introduces a framework for quantifying classifier utility under LDP mechanisms. The framework connects the concentration analysis of LDP mechanisms with the robustness analysis of classifiers. It provides guidelines for selecting the best mechanism and privacy parameter for a given classifier. Case studies on typical classifiers under LDP mechanisms demonstrate the framework’s effectiveness and applicability.
@@ 8 Conclusions
Ethical Considerations
This paper does not involve any human subjects or animals. While we use real-world medical and financial data, they are publicly available, and all personal information has been anonymized.
Open Science
All code for the quantification framework, evaluations, and datasets used in this paper will be made publicly available upon publication. We will include scripts for reproducing the experiments and generating the figures presented in the paper. Meanwhile, we provide an anonymized code snapshot for review that is currently available at https://anonymous.4open.science/r/BoundClassifier-LDP-USENIX/README.md.
@@ Appendix A Related Work
A.1 Closely Related Concepts
We first discuss three closely related concepts and their differences to our work.
@@ A.1 Closely Related Concepts
A number of the existing literature focuses on empirical evaluations of data utility under semantic perturbations [47], i.e. perturbations that involve semantic meanings. For image classifiers, such perturbations include brightness adjustments [6], blur [13], and data augmentation such as rotation and scaling [32, 49]. For text classification, examples include synonym replacement and random swap [58, 34], etc. For speech recognition, there are acoustic distortions [36]. These perturbations do not provide formal privacy guarantees, and some of them even preserve the semantic meanings of the original data.
This paper focuses on a specific category of mathematically tractable perturbations, LDP mechanisms. These perturbations, such as the Laplace or Gaussian noise, have weaker semantic meanings but provide provable privacy guarantees against adversaries.
@@
The term concentration property mostly relates to the concentration inequalities in probability theory, e.g. Chernoff bound and Hoeffding inequality (see book [14]). These inequalities provide upper bounds on the probability that the sum of independent random variables deviates from its expected value. From these inequalities, the notion $(\alpha,\beta)$-accuracy [14, 10] is defined as: the error is bounded by $\alpha$ with probability at least $\beta$. In DP, this notion has been used to analyze the error of the summation query in the shuffle model [10]. It also appears in accuracy analysis of data exploring languages under DP guarantee, such as Haskell and SQL [53, 22].
@@
Robustness refers to a classifier’s ability to maintain performance (utility) under perturbations [18]. This property is crucial as classifiers often encounter data that deviates from their training distribution. Robustness also serves as a metric for evaluating a classifier’s generalization ability [65]. Beyond semantic perturbations, robustness has been extensively studied under $\ell_{p}$-norm perturbations [18, 40, 39], where input data is perturbed within an $\ell_{p}$-norm ball.
@@
A.2 More Broad Related Work
@@
Fairness dictates that two individuals who are similar w.r.t. a particular task should be classified similarly. Under some metrics of similarity, it relates to the concept of DP [16]. DP mechanisms can impact the fairness of classifiers by changing biases in the input data distribution. For instance, if certain groups are over- or under-represented in the data shared with classifiers, it can lead to biased predictions after the application of DP mechanisms [41, 3].
@@
This paper addresses data privacy when (L)DP mechanisms are applied by users before sharing data with classifiers, i.e. ensuring privacy during classifier usage. An orthogonal line of research examines privacy risks posed by classifiers themselves, focusing on data leakage during training. Such works aim to counter membership inference attacks [20, 51], which infer training data from model parameters. DP-SGD [1] is a well-known defense, applying Gaussian noise to gradient updates during training.
Federated learning [43] is a stricter privacy setting, where sensitive training data is distributed across multiple users and cannot be shared with a central server. In this context, DP-Fed-SGD [44] perturbs local gradients before aggregation to the central server. For a comprehensive overview, see a recent survey [21].
@@
Data utility is the most crucial metric for LDP mechanisms. For simple queries such as mean, summation, and histogram [55, 54, 56], data utility is essentially determined by the theoretical mean squared error (MSE) of the mechanism. While empirical evaluations can be conducted by sampling from the mechanism, the empirical utility will converge to the theoretical MSE when the sample size is large enough. Meanwhile, MSE is also the most common metric in mechanism design and analysis.
However, a mechanism with lower MSE does not always guarantee better utility for all applications, particularly classifiers. This paper has shown that classifier utility depends on both the concentration property of the mechanism and the classifier’s robustness. We provide a framework to quantify classifier utility under LDP mechanisms, offering an alternative to empirical evaluations.
@@ B.1 Proof of Theorem 3
B.2 Proof of Theorem 4 and Discussion
B.2.1 Our Gaussian Mechanism
The proof generally follows the structure of Dwork’s proof of the Gaussian mechanism on page 261 [17]. Their proof assumes $\varepsilon\leq 1$ for the soundness of the $\ln(\varepsilon,\Delta f)$ function. We eliminate this limitation by refining the proof technique and providing a new noise bound.
@@ B.2.1 Our Gaussian Mechanism
For the Gaussian distribution and an original data $x_{1}\in[0,1]$, the probability of seeing $x_{1}+\tilde{x}$ with $\tilde{x}\sim\mathcal{N}(0,\sigma^{2})$ is given by the probability density function at $\tilde{x}$, i.e. $pdf[\tilde{x}$]. The probabilities of seeing this same value when the original data is $x_{2}\in[0,1]$ belongs to a set $\{pdf[\tilde{x}+\Delta]:\Delta\in[-1,1]\}$. Thus, the privacy loss is
$\begin{split}\mathcal{L}_{\mathcal{M},x_{1},x_{2}}(\tilde{x})&=\ln\left(\frac{pdf[\tilde{x}]}{pdf[\tilde{x}+\Delta]}\right)=\ln\frac{\exp(-\frac{1}{2\sigma^{2}}\cdot\tilde{x}^{2})}{\exp(-\frac{1}{2\sigma^{2}}\cdot(\tilde{x}+\Delta)^{2})}\\ &=\frac{1}{2\sigma^{2}}\cdot\left(2\tilde{x}\Delta+\Delta^{2}\right).\end{split}$
@@
B.2.2 Comparison with Dwork’s Proof.
@@
Specifically, (i) the proof is based on the assumption that the privacy loss can be probabilistically bounded by $1-\delta$, which is consistent with the PAC privacy notion but hard to interpret in the original $(\varepsilon,\delta)$ privacy notion. The $(\varepsilon,\delta)$ privacy notion says that the distance between two distributions is always bounded by the given $(\varepsilon,\delta)$ pair, i.e. strictly cannot be unbounded. From this notion, it is hard to interpret the probabilistic bound of $1-\delta$ in Dwork’s proof. (ii) The proof is based on a tail bound
@@
where $Z\sim\mathcal{N}(0,\sigma^{2})$. This bound is problematic. As a counterexample, if $t=0$ and $\sigma=1$, the bound gives $\Pr[Z\geq 0]\leq 1/\sqrt{2\pi}\approx 0.399$, which is incorrect since $\Pr[Z\geq 0]=0.5$ for Gaussian distributions with mean 0.
B.2.3 Comparison with the Analytic Gaussian.
Another Gaussian mechanism is the analytic Gaussian mechanism [5]. The name originates from the analytic form of the Gaussian distribution in its result. We don’t adopt this mechanism for two reasons: (i) It lacks analytical form of noise scale $\sigma$ due to the complexity of the given formula to satisfy their privacy definition. i.e. Theorem 8 in [5]. As stated in their paper: "we propose to find $\sigma$ using a numerical algorithm… (p4, line 7)", which complicates analytical utility quantification. (ii) It is based on an alternative privacy definition:
@@ B.3 Proof of Theorem 2
The combination of $d$ independent $(\varepsilon,\delta)$-PAC LDP mechanisms satisfies pure $\varepsilon$-LDP only if all $d$ mechanisms satisfy pure $\varepsilon$-LDP. Otherwise, the failure probability is at least $\delta$. Therefore, the probability of no failure is at least $(1-\delta)^{d}$. Thus, the total failure probability of $d$ independent $(\varepsilon,\delta)$-PAC LDP mechanisms is $1-(1-\delta)^{d}$.
@@
Remark. Compare with the original combination theorem for $d$ mechanisms, i.e. $(d\varepsilon,d\delta)$-LDP [17], our result is guaranteed to be more precise, as $1-(1-\delta)^{d}\leq d\delta$ for any $\delta\in(0,1)$ and $d\geq 1$.
@@ Appendix C Complementary Materials
C.1 Deterministic Robustness of White-box Classifiers (Section 4.2)
@@
C.1.1 Closed-form Classifiers
A closed-form classifier has a mathematical expression that can be directly analyzed. Examples include the naive Bayes classifier [59], QDA [24, 4], certain variants of SVM [28], and clustering models like $k$-means [61] and Gaussian mixture models [27]. Closed-form expression means excellent interpretability, including robustness analysis. For these classifiers, decision boundaries can be analytically derived from the model parameters, allowing for analytical expression of robustness radius $\theta$ for any input variable $x_{0}$.
@@ C.1.1 Closed-form Classifiers
C.1.2 Non-closed-form Classifiers
Non-closed-form classifiers often lack explicit mathematical expressions for their decision boundaries. Examples include neural network [60] and decision tree [8]. Their decision boundaries are analytically intractable, making it challenging to derive the robustness radius $\theta$ directly. Finding $\theta$ for these classifiers is known as the robustness verification problem [33, 45], i.e. verify whether $h(B_{\theta}(x))=h(x)$ for a given $\theta$ at $x$. The main insight is to encode classifiers as a set of constraints, transforming the problem of verifying $h(B_{\theta}(x))=h(x)$ into a satisfiability problem that can be solved by constraints solvers.

### v3

@@ 7.2.2. Bank Customer Attrition
Figure 7. Empirical and theoretical utility for two classifiers trained on the Bank Customer Attrition dataset.
@@
7.2.3. MNIST-$\times7$
Improvements for high-dimensional classifiers. (i) In practice, not all dimensions are sensitive and require privacy protection. While we treat all dimensions as sensitive in the MNIST-$\times7$ dataset, users can opt to apply LDP mechanisms to only a subset of dimensions, i.e. specific parts of the image. In such cases, the theoretical utility quantification can be more accurate. (ii) At present, we identify a connected (local) robustness rectangle that contains $x$; however, the classifier may admit additional disconnected and non-rectangular robust regions. One can use Monte Carlo sampling to discover such regions and then combine this with Monte Carlo integration to obtain a more accurate theoretical utility quantification for high-dimensional classifiers. Appendix C.8.1 provides details.
@@ 7.2.3. MNIST-$\times7$
Figure 8. Empirical and theoretical utility for a Neural Network classifier trained on the MNIST-$\times7$ dataset.
Average-case and worst-case analysis. We use the first 50 images from the MNIST-$\times7$ dataset as representative samples for analyzing the average-case and worst-case utility under LDP-perturbed inputs. Appendix C.8.2 presents the results of this analysis.
7.3. Summary of Case Studies
We presented case studies of theoretical utility statements for different classifiers under inputs perturbed by typical LDP mechanisms and compared them with empirical utility. The key findings from our case studies are as follows:
@@ 7.3. Summary of Case Studies
8. Conclusions
This paper introduces a framework for quantifying classifier utility under LDP-perturbed inputs. The framework connects the concentration analysis of LDP mechanisms with the robustness analysis of classifiers. It provides guidelines for selecting the best mechanism and privacy parameter for a given classifier. Beyond the core framework, we introduce two novel refinement techniques that further improve utility quantification. Case studies on typical classifiers under LDP-perturbed inputs demonstrate the framework’s effectiveness and applicability.
@@ 8. Conclusions
Acknowledgments
We thank the anonymous reviewers and the revision editor for their valuable feedback and guidance, which significantly improved this paper. We also acknowledge the use of GPT-5 for language refinement in this paper.
@@ Appendix A Related Work
This paper studies the data utility of classifiers under perturbed inputs. We first review closely related concepts—semantic perturbations, concentration properties of random variables, and classifier robustness—and clarify how they differ from our setting. We then discuss broader related work.
A.1. Closely Related Concepts
@@
A number of the existing literature focuses on empirical evaluations of data utility under semantic perturbations (DBLP:conf/cvpr/MohapatraWC0D20), i.e. perturbations that involve semantic meanings. For image classifiers, such perturbations include brightness adjustments (bendjillali:hal-03564174), blur (DBLP:conf/qomex/DodgeK16), and data augmentation such as rotation and scaling (DBLP:conf/cvpr/KanbakMF18; DBLP:journals/corr/abs-1712-04621). For text classification, examples include synonym replacement and random swap (DBLP:conf/emnlp/WeiZ19; DBLP:conf/naacl/Kobayashi18), etc. For speech recognition, there are acoustic distortions (DBLP:journals/taslp/LiDGH14). These perturbations do not provide formal privacy guarantees, and some of them even preserve the semantic meanings of the original data.
This paper focuses on a specific category of mathematically tractable perturbations, LDP mechanisms. These perturbations, such as the Laplace or Gaussian noise, have weaker semantic meanings but provide provable privacy guarantees against adversaries. Additionally, there are variants of LDP mechanisms tailored for semantic perturbations, such as sentence embedding perturbation (DBLP:conf/www/DuYCS23). The LDP mechanisms examined in this paper are typical building blocks for numerous LDP protocols in practical applications.
@@
The term concentration property mostly relates to the concentration inequalities in probability theory, e.g. Chernoff bound and Hoeffding inequality (see book (dubhashi2009concentration)). These inequalities provide upper bounds on the probability that the sum of independent random variables deviates from its expected value. From these inequalities, the notion $(\alpha,\beta)$-accuracy (dubhashi2009concentration; DBLP:conf/eurocrypt/CheuSUZZ19) is defined as: the error is bounded by $\alpha$ with probability at least $\beta$. In DP, this notion has been used to analyze the error of the summation query in the shuffle model (DBLP:conf/eurocrypt/CheuSUZZ19). It also appears in accuracy analysis of data exploring languages under DP guarantee, such as Haskell and SQL (DBLP:conf/sp/VesgaRG20; DBLP:conf/sigmod/0002HIM19).
@@
Robustness refers to a classifier’s ability to maintain performance (utility) under input perturbations (DBLP:conf/nips/FawziMF16). This property is crucial as classifiers often encounter data that deviates from their training distribution. Robustness also serves as a metric for evaluating a classifier’s generalization ability (DBLP:journals/pami/ZhouLQXL23). Beyond semantic perturbations, robustness has been extensively studied under $\ell_{p}$-norm perturbations (DBLP:conf/nips/FawziMF16; DBLP:conf/iclr/MadryMSTV18; DBLP:journals/csur/MachadoSG23), where input data is perturbed within an $\ell_{p}$-norm ball. Within this perturbation model, various verification techniques have been developed to analyze classifier robustness. In addition to approximating the robustness of a classifier as described in Section 4.2, we can also approximate the classifier using a “smoothed” classifier via randomized smoothing (pmlr-v54-mcmahan17a) and then analyze the robustness radius of the smoothed classifier.
@@
A.2. Broader Related Work
@@
Fairness dictates that two individuals who are similar w.r.t. a particular task should be classified similarly. Under some metrics of similarity, it relates to the concept of DP (DBLP:conf/innovations/DworkHPRZ12). DP mechanisms can impact the fairness of classifiers by changing biases in the input data distribution. For instance, if certain groups are over- or under-represented in the data provided to a classifier, applying DP mechanisms can shift the input distribution and thereby alter the fairness (or bias) of the resulting predictions (DBLP:conf/csfw/MakhloufSAP24; DBLP:conf/dbsec/ArcoleziMP23).
@@
This paper addresses data privacy when (L)DP mechanisms are applied by users before sharing data with classifiers, i.e. ensuring privacy during classifier usage. An orthogonal line of research examines privacy risks posed by classifiers themselves, focusing on data leakage during training. Such works aim to counter membership inference attacks (DBLP:conf/uss/FredriksonLJLPR14; DBLP:conf/sp/ShokriSSS17), which infer training data from model parameters. DP-SGD (DBLP:conf/ccs/AbadiCGMMT016) is a well-known defense, applying Gaussian noise to gradient updates during training.
Federated learning (pmlr-v54-mcmahan17a) is a stricter privacy setting, where sensitive training data is distributed across multiple users and cannot be shared with a central server. In this context, DP-Fed-SGD (brendan2018learning) perturbs local gradients before aggregation to the central server. For a comprehensive overview, see a recent survey (DBLP:journals/corr/abs-2405-08299).
@@
Data utility is the most crucial metric for LDP mechanisms. For simple queries such as mean, summation, and histogram (DBLP:conf/uss/WangBLJ17; DBLP:conf/icde/WangXYZHSS019; DBLP:conf/sp/WangLJ18), data utility is essentially determined by the theoretical mean squared error (MSE) of the mechanism. While empirical evaluations can be conducted by sampling from the mechanism, the empirical utility will converge to the theoretical MSE when the sample size is large enough. Meanwhile, MSE is also the most common metric in mechanism design and analysis.
However, a mechanism with lower MSE does not always guarantee better utility for all applications, particularly classifiers. This paper has shown that classifier utility depends on both the concentration property of the mechanism and the classifier’s robustness. We provide a framework to quantify classifier utility under LDP-perturbed data, offering an alternative to empirical evaluations.
@@ B.1. Proof of Theorem 5
B.2. Proof of Theorem 6 and Discussion
B.2.1. Our Gaussian Mechanism
The proof generally follows the structure of Dwork’s proof of the Gaussian mechanism on page 261 in (DBLP:journals/fttcs/DworkR14). Their proof assumes $\varepsilon\leq 1$ for the soundness of the $\ln(\varepsilon,\Delta f)$ function. We eliminate this limitation by refining the proof technique and providing a new noise bound.
@@ B.2.1. Our Gaussian Mechanism
For the Gaussian distribution and an original data $x_{1}\in[0,1]$, the probability of seeing $x_{1}+\tilde{x}$ with $\tilde{x}\sim\mathcal{N}(0,\sigma^{2})$ is given by the probability density function at $\tilde{x}$, i.e. $pdf[\tilde{x}$]. The probabilities of observing the same perturbed value when the original input is $x_{2}\in[0,1]$ form the set $\{\,pdf[\tilde{x}+\Delta]:\Delta\in[-1,1]\,\}$. Thus, the privacy loss is
$\begin{split}\mathcal{L}_{\mathcal{M},x_{1},x_{2}}(\tilde{x})&=\ln\left(\frac{pdf[\tilde{x}]}{pdf[\tilde{x}+\Delta]}\right)=\ln\frac{\exp(-\frac{1}{2\sigma^{2}}\cdot\tilde{x}^{2})}{\exp(-\frac{1}{2\sigma^{2}}\cdot(\tilde{x}+\Delta)^{2})}\\ &=\frac{1}{2\sigma^{2}}\cdot\left(2\tilde{x}\Delta+\Delta^{2}\right)=\frac{\Delta}{\sigma^{2}}\cdot\tilde{x}+\frac{\Delta^{2}}{2\sigma^{2}}.\end{split}$
@@
B.2.2. Comparison with Dwork’s Proof.
@@
Specifically, (i) the proof is based on the assumption that the privacy loss can be probabilistically bounded by $1-\delta$, which is consistent with the PAC privacy notion but hard to interpret in the original $(\varepsilon,\delta)$ privacy notion. The $(\varepsilon,\delta)$ privacy notion intuitively says that the distance between two distributions is always bounded by the given $(\varepsilon,\delta)$ pair, i.e. strictly cannot be unbounded. From this notion, it is hard to interpret the probabilistic bound of $1-\delta$ in Dwork’s proof. (ii) The proof is based on a tail bound (p262 in (DBLP:journals/fttcs/DworkR14)):
@@
where $Z\sim\mathcal{N}(0,\sigma^{2})$. This bound is problematic. As a counterexample, if $t=0$ and $\sigma=1$, the bound gives $\Pr[Z\geq 0]\leq 1/\sqrt{2\pi}\approx 0.399$, which is incorrect since $\Pr[Z\geq 0]=0.5$ for arbitrary Gaussian distributions with mean 0.
B.2.3. Comparison with the Analytic Gaussian.
Another Gaussian mechanism is the analytic Gaussian mechanism (DBLP:conf/icml/BalleW18). The name originates from the analytic form of the Gaussian distribution in its result. We don’t adopt this mechanism for two reasons: (i) It lacks analytical form of noise scale $\sigma$ due to the complexity of the given formula to satisfy their privacy definition. i.e. Theorem 8 in (DBLP:conf/icml/BalleW18). As stated in their paper: “we propose to find $\sigma$ using a numerical algorithm… (p4, line 7)”, which complicates analytical utility quantification. (ii) It is based on an alternative privacy definition:
@@ B.3. Proof of Theorem 4
We calculate the total failure probability of $d$ mechanisms to show that it is $1-(1-\delta)^{d}$.
The combination of $d$ independent $(\varepsilon,\delta)$-PAC LDP mechanisms satisfies pure $\varepsilon$-LDP only if all $d$ mechanisms satisfy pure $\varepsilon$-LDP. Therefore, the probability of no failure is at least $(1-\delta)^{d}$ by independence. Thus, the total failure probability of $d$ independent $(\varepsilon,\delta)$-PAC LDP mechanisms is at most $1-(1-\delta)^{d}$.
@@
Remark. Compare with the original combination theorem for $d$ mechanisms, i.e. $(d\varepsilon,d\delta)$-LDP (DBLP:journals/fttcs/DworkR14), our result is guaranteed to be more precise, as $1-(1-\delta)^{d}\leq d\delta$ for any $\delta\in(0,1)$ and $d\geq 1$.
@@ Appendix C Complementary Materials
Notations used in this paper are summarized in Table 3.
Table 3. Notations used in this paper.
Notation Meaning
$\mathcal{M}(x)$ LDP mechanism applied to input $x$
$\mathcal{X}$ Input domain of an LDP mechanism
$\tilde{x}=\mathcal{M}(x)$ Perturbed version of $x$ (sent to $h$)
$h$ Classifier function
$B_{\theta}(x)$ $\theta$-ball centered at $x$
$\rho(\varepsilon,\theta)$ Utility quantification of a classifier
$F_{\mathcal{M}}(\cdot)$ CDF of LDP mechanism $\mathcal{M}$
$\omega$ $1-\omega$ indicates the confidence level
$\tau$ Robustness tolerance
C.1. Deterministic Robustness of White-box Classifiers (Section 4.2)
@@
C.1.1. Closed-form Classifiers
A closed-form classifier has a mathematical expression that can be directly analyzed. Examples include the naive Bayes classifier (naive_bayes), QDA (DBLP:journals/corr/abs-1906-02590; lda_qda), certain variants of SVM (DBLP:books/lib/HastieTF09), and clustering models like $k$-means (wiki_k-means) and Gaussian mixture models (GMM). Closed-form expression means excellent interpretability, including robustness analysis. For these classifiers, decision boundaries can be analytically derived from the model parameters, allowing for analytical expression of robustness radius $\theta$ for any input variable $x_{0}$.
@@ C.1.1. Closed-form Classifiers
C.1.2. Non-closed-form Classifiers
Non-closed-form classifiers often lack explicit mathematical expressions for their decision boundaries. Examples include neural network (DBLP:conf/iclr/WengZCYSGHD18) and decision tree (DBLP:conf/nips/ChenZS0BH19). Their decision boundaries are analytically intractable, making it challenging to derive the robustness radius $\theta$ directly. Finding $\theta$ for these classifiers is known as the robustness verification problem (DBLP:conf/cav/KatzBDJK17; DBLP:journals/corr/abs-2206-12227), i.e. verify whether $h(B_{\theta}(x))=h(x)$ for a given $\theta$ at $x$. The main insight is to encode classifiers as a set of constraints, transforming the problem of verifying $h(B_{\theta}(x))=h(x)$ into a satisfiability problem that can be solved by constraints solvers.